OpenStudy (anonymous):

Jim needs 3 1/2 yards of burlap to cover his bulletin board. His brother gives him 1 1/6 yards. How much more does Jim need to purchase?

OpenStudy (shrutipande9):

try posting this in maths subject

OpenStudy (shrutipande9):

but u need to do....\[3\frac{ 1 }{ 2 } - 1\frac{ 1 }{ 6 }\]

OpenStudy (shrutipande9):

\[\frac{ (3x2) + 1 }{ 2 } - \frac{ (6x1) + 1 }{ 6 }\]